OpenStudy (anonymous):

Proving Identities: sin^3(x)+cos^3(x)+sin(x)cos^2(x)+sin^2(x)cos(x) = sin(x)+cos(x) ? how?

OpenStudy (psymon):

Ooo monster. *looks*

OpenStudy (anonymous):

maybe factor?

OpenStudy (anonymous):

not equal? the book says prove the identities, it didn't say anything about it being an identity or not. (_ _)"

OpenStudy (psymon):

Be kind of mean if the book gave you something that could not be proven, lol.

OpenStudy (anonymous):

There must be a typo in it somewhere o.O

OpenStudy (anonymous):

hmmm.

OpenStudy (anonymous):

oh wait. there is.

OpenStudy (psymon):

Well, I didn't prove it BUT I did give up and just turn everything into a ton ton ton of sinx's.

OpenStudy (anonymous):

lol. I was about to start manipulation, but I thought I would check just to make sure it was true.

OpenStudy (anonymous):

i'm sorry. the last term on the left hand side, the cosx doesn't have to be squared.

OpenStudy (psymon):

so just sin^2(x)cosx?

OpenStudy (anonymous):

yes sir.

OpenStudy (psymon):

Oh, well that one turns out just fine then :D

OpenStudy (anonymous):

:) sorry for my mistake. so... how it can be proved?

OpenStudy (psymon):

Do this with the first two terms. \[(\sin ^{2}x)(sinx) + (\cos ^{2}x)(cosx)\] Now use the pythagorean identity on the sin^2(x) and the cos^2(x) in order to change the sin^2(x) into cosines and the cos^2(x) into sines

OpenStudy (psymon):

Use \[\sin ^{2}x + \cos ^{2}x = 1\] And solve for sin^2(x) in thefirst term and solve for cos^2(x) in the 2nd term. Then things will multiply and cancel out stuffz.

OpenStudy (anonymous):

ok. (solving on paper)

OpenStudy (anonymous):

done. :) thanks sir.

OpenStudy (anonymous):

You could also use the fact that:\[s^3+c^3=(s+c)(s^2-sc+c^2)=(s+c)(1-sc)\]and:\[s^2c+sc^2=sc(s+c).\]Im using s for sin and c for cos because im lazy :P

OpenStudy (psymon):

See, you could've done that, but we ain't fancy folk D:

OpenStudy (anonymous):

lolol.

OpenStudy (anonymous):

(_ _) sorry because in my notebook both sin and cos are squared on the last term in the left hand side. but when I look at the book there goes the truth. sorry for disturbing you.

OpenStudy (psymon):

Textbooks make mistakes more often than we'd like.